OpenStudy (anonymous):

Factor and simplify (cos^2(x)-1)/(cos(x) +1)

OpenStudy (anonymous):

Distribute using the FOIL method, then combine like terms.

OpenStudy (anonymous):

let x = cosx than x^2 -1/ x +1 factor x^2 -1 = (x+1)(x+1)/(x+1) = x-1 so the solution for x = 1 so solve for cosx =1

OpenStudy (anonymous):

\[\frac{x^2-1}{x+1}=\frac{(x+1)(x-1)}{x+1}=x-1\]

OpenStudy (anonymous):

Thanks